OpenStudy (anonymous):

how does 150cm^2 = 23.2in^2

OpenStudy (anonymous):

1inch = 2.54cm 1cm=1/2.54 1cm^2=1/(2.54*2.54) 150cm^2=150*1/(2.54*2.54) 150cm^2=23.2in^2

OpenStudy (anonymous):

in^2 = 150 cm^2 * (1 in)^2/ (2.54cm)^2 = 23.25 in^2